# On the Unknowable Limits to Prediction

### v3

@@ 3. Predictive Model
Substituting for Equations S2-S4, we obtain:
@@
It then logically follows that predicting the target variable as accurately as can be done in practice (see Section SI.1.1) results in the case where all epistemic errors are eliminated, and only aleatoric error remains: $\text{Error}=y_{\text{pred}}-y_{\text{true}}=\varepsilon$.
SI.1.4 Relationship to Existing Frameworks
Our framework mainly decomposes errors in a conceptually simple way (i.e., up until now, we have been dealing with scalar target outcome observations). We next develop the aggregate level decomposition of the expected sum of squared errors and relate our framework to the bias, variance, and error frameworks of canonical textbooks (e.g., james2013introduction; hastie2009elements) as well as papers (pedro2000unified; james2003variance) in the field of statistical learning. Suppose we have a sample with $N$ observations, for each $y_{\text{true},i}\in\textbf{y}_{\text{true}}$ and $y_{\text{pred},i}\in\textbf{y}_{\text{predicted}}$ for $i=\{1,2,...,N\}$. Equation LABEL:eq:error_decomposition holds as follows:
@@
$=\delta^{\text{error}}_{\mathbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon,$ (S10)
@@
  • $\delta^{\text{error}}_{\textbf{x}}\coloneqq{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})$; the difference in predicted values owing to $\delta_{\textbf{x}}$. $\mathbb{E}[\delta^{\text{error}}_{\textbf{x}}]=\mu^{\text{error}}_{\textbf{x}}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})=\sigma^{2}_{\delta^{\text{error}}_{\textbf{x}}}$. As $\delta^{\text{error}}_{\textbf{x}}$ is expressed as the difference between the fitted values, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})$ depends on the specific form of $f$, the distribution of $\textbf{x}_{\text{observed}}\text{ and }\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}$, and their dependencies.
  • $\delta^{\text{error}}_{y}\coloneqq{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})$; the difference in predicted values owing to $\delta_{y}$. $\mathbb{E}[\delta^{\text{error}}_{y}]=\mu^{\text{error}}_{y}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{y})=\sigma^{2}_{\delta^{\text{error}}_{y}}$. Similarly, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{y})$ depends on the specific form of $f$, the distribution of $y_{\text{observed}}\text{ and }y_{\text{observed}}-\delta_{y}$, and their dependencies.
  • $\delta_{f}$ is defined as before in Equation S6.
@@
$=\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{f}.$ (S11)
@@
$=\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]+\operatorname{Var}(\varepsilon).$ (S12)
With Equations S11-S12 we can write our squared error loss as:
@@
$=(\underbrace{\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{\delta_{f}}}_{\text{Bias}})^{2}+\underbrace{\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]}_{\text{Variance}}+\underbrace{\operatorname{Var}(\varepsilon)}_{\text{AleatoricError}},$ (S13)
@@
which results in the familiar Irreducible $\text{Irreducible Error}+\text{Bias}^{2}+\text{Variance}$ framework akin to Equation 7.9 of hastie2009elements. For simplicity, we don’t expand $\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]$, of which the full expansion contains the covariance terms between each of the $\delta$ terms.
SI.1.5 Partially captured aleatoric error in observed data
@@
In Equations S3-S4, we assume that $y_{\text{observed}}$ captures all aleatoric error in $y_{\text{true}}$. However, since $y_{\text{observed}}$ is measured after constructing $y_{\text{true}}$, this assumption may not hold in all situations. For example, observed data may be a biased sub-sample of the whole population. With such observed data, we can only optimize epistemic errors ($\delta_{f},\delta_{\textbf{x}}\text{ and }\delta_{y}$) that are inherent in training data. In this case – even when all epistemic error is reduced – $\varepsilon_{y_{\text{observed}}}$ cannot represent the true $\varepsilon_{y}$. Consider a true target variable which is measurable across a population ($y_{\text{true}}$), but $y_{\text{observed}}$ only covers a non-representative 70% of the population. In this case, $\varepsilon_{y_{\text{observed}}}$ only captures the $\varepsilon_{y}$ of the observed sample, and lacks the ability to capture the component which would be inherent in the remaining 30% of the population. The representativeness of the error term $\varepsilon_{y_{\text{observed}}}$ depends on how well $y_{\text{observed}}$ represents true reality ($y_{\text{true}}$). It also affects assumptions we need to make in terms of the bias-variance decomposition as per Equations S12-S13. If $\varepsilon_{y_{\text{observed}}}\neq\varepsilon_{y}$, the covariance between $\varepsilon_{y_{\text{observed}}}$ and $\delta_{f},\delta^{\text{error}}_{y}\text{ and }\delta^{\text{error}}_{\textbf{x}}$ are no longer zero. The $\operatorname{Var}(Error)$ term must also incorporate $\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta_{f}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{y}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{\textbf{x}})$. Therefore, guaranteeing the representativeness of $y_{\text{observed}}$ to $y_{\text{true}}$ (which also affects the representativeness of $\textbf{x}_{\text{observed}}$ to $\textbf{x}_{\text{true}}$ from a measurement perspective) is a necessary step before considering epistemic errors in the form of $\delta_{f},\delta_{y},\text{ and }\delta_{\textbf{x}}$, as it interrogates whether the questions we are answering are really what we meant to ask.
@@
Figure S1: Decomposed Learning Curves: Two complimentary example. Panel ‘a.’ represents a target variable with relatively little aleatoric error, and a rapid learning rate with large gains from eliminating measurement error. Panel ‘b.’ represents a target variable with relatively higher aleatoric error, a slower learning rate, and relatively fewer gains from improving measurement. Note: there is no requirement for such learning curves to be continuous functions. Dashed horizontal red lines denote what is commonly known as the ‘predictive ceiling’ (no prediction error). Golden arrows denote the rate at which reducible error is reducing. Dashed blue lines represents a less accurate learning algorithm, and the solid blue line represents the case where learning error is entirely eliminated.
@@
  • James et al. (2013) James, G., D. Witten, T. Hastie, R. Tibshirani, et al. (2013). An introduction to statistical learning Volume 112. Springer.
  • Pedro (2000) Pedro, D. (2000). A unified bias-variance decomposition and its applications. In 17th International conference on machine learning, pp.  231–238.

### v4

@@ 1.1.3 Error Decomposition
Substituting for Supplementary Information Equations 1-3, we obtain:
@@
It then logically follows that predicting the target variable as accurately as can be done in practice (see Supplementary Information Section 1.1.1) results in the case where all epistemic errors are eliminated, and only aleatoric error remains: $\text{Error}=y_{\text{pred}}-y_{\text{true}}=\varepsilon$.
1.1.4 Relationship to Existing Frameworks
Our framework mainly decomposes errors in a conceptually simple way (i.e., up until now, we have been dealing with single scalar target outcome observations). We next develop the aggregate level decomposition of the expected sum of squared errors and relate our framework to the bias, variance, and error frameworks of canonical textbooks (e.g., james2013introduction; hastie2009elements) as well as papers (pedro2000unified; james2003variance) in the field of statistical learning. Suppose we have a sample with $N$ observations, with each $y_{\text{true},i}\in\textbf{y}_{\text{true}}$ and $y_{\text{pred},i}\in\textbf{y}_{\text{predicted}}$ for $i=\{1,2,...,N\}$. Supplementary Information Equation LABEL:eq:error_decomposition holds as follows:
@@ 1.1.4 Relationship to Existing Frameworks
$=\delta^{\text{error}}_{\mathbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon,$ (9)
@@
  • $\delta^{\text{error}}_{\textbf{x}}\coloneqq{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})$; the difference in predicted values owing to $\delta_{\textbf{x}}$. $\mathbb{E}[\delta^{\text{error}}_{\textbf{x}}]=\mu^{\text{error}}_{\textbf{x}}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})=\sigma^{2}_{\delta^{\text{error}}_{\textbf{x}}}$. As $\delta^{\text{error}}_{\textbf{x}}$ is expressed as the difference between the fitted values, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})$ depends on the specific form of $f$, the distribution of $\textbf{x}_{\text{observed}}\text{ and }\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}$, and their dependencies;
  • $\delta^{\text{error}}_{y}\coloneqq{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})$; the difference in predicted values owing to $\delta_{y}$. $\mathbb{E}[\delta^{\text{error}}_{y}]=\mu^{\text{error}}_{y}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{y})=\sigma^{2}_{\delta^{\text{error}}_{y}}$. Similarly, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{y})$ depends on the specific form of $f$, the distribution of $y_{\text{observed}}\text{ and }y_{\text{observed}}-\delta_{y}$, and their dependencies;
  • $\delta_{f}$ is defined as before in Equation 5.
@@
$=\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{f}+\varepsilon.$ (10)
@@
$=\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]+\operatorname{Var}(\varepsilon).$ (11)
With Supplementary Information Equations 10-11, we can write our squared error loss as:
@@
$=(\underbrace{\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{\delta_{f}}+\varepsilon}_{\text{Bias}})^{2}+\underbrace{\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})+]}_{\text{Variance}}+\underbrace{\operatorname{Var}(\varepsilon)}_{\text{Aleatoric Error}},$ (12)
@@
which results in the familiar Irreducible $\text{Irreducible Error}+\text{Bias}^{2}+\text{Variance}$ framework akin to Equation 7.9 of hastie2009elements. For simplicity, we don’t expand $\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]$, which would include the covariance between each of the $\delta$ terms.
1.1.5 Partially captured aleatoric error in observed data
@@
In Supplementary Information Equations 2-3, we assume that $y_{\text{observed}}$ captures all aleatoric error in $y_{\text{true}}$. However, since $y_{\text{observed}}$ is measured after constructing $y_{\text{true}}$, this assumption may not hold in all situations. For example, observed data may be a biased sub-sample of the whole population. With such observed data, we can only optimize epistemic errors ($\delta_{f},\delta_{\textbf{x}}\text{ and }\delta_{y}$) that are inherent in training data. In this case – even when all epistemic error is reduced – $\varepsilon_{y_{\text{observed}}}$ cannot represent the true $\varepsilon_{y}$. Consider a true target variable which is measurable across a population ($y_{\text{true}}$), but $y_{\text{observed}}$ only covers a non-representative 70% of the population. In this case, $\varepsilon_{y_{\text{observed}}}$ only captures the $\varepsilon_{y}$ of the observed sample, and lacks the ability to capture the component which would be inherent in the remaining 30% of the population. The representativeness of the error term $\varepsilon_{y_{\text{observed}}}$ depends on how well $y_{\text{observed}}$ represents true reality ($y_{\text{true}}$). It also affects assumptions we need to make in terms of the bias-variance decomposition as per Equations 11-12. If $\varepsilon_{y_{\text{observed}}}\neq\varepsilon_{y}$, the covariances between $\varepsilon_{y_{\text{observed}}}$ and $\delta_{f},\delta^{\text{error}}_{y}\text{ and }\delta^{\text{error}}_{\textbf{x}}$ are no longer zero. The $\operatorname{Var}(Error)$ term must also incorporate $\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta_{f}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{y}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{\textbf{x}})$. Therefore, guaranteeing the representativeness of $y_{\text{observed}}$ to $y_{\text{true}}$ (which also affects the representativeness of $\textbf{x}_{\text{observed}}$ to $\textbf{x}_{\text{true}}$ from a measurement perspective) is a necessary step before considering epistemic errors in the form of $\delta_{f},\delta_{y},\text{ and }\delta_{\textbf{x}}$, as it interrogates whether the questions we are answering are really what we meant to ask.
@@
Figure 1: Supplementary Figure 1: Additional Learning Curve Examples. Two complementary examples to Figure 1. Panel ‘a.’ represents a target variable with relatively little aleatoric error, and a rapid learning rate with large gains from eliminating measurement error. Panel ‘b.’ represents a target variable with relatively higher aleatoric error, a slower learning rate, and relatively fewer gains from improving measurement. Note: there is no requirement for such learning curves to be continuous functions. Dashed horizontal red lines denote what is commonly known as the ‘predictive ceiling’ (no prediction error). Golden arrows denote the rate at which reducible error is reducing. Dashed blue lines represents a less accurate learning algorithm, and the solid blue line represents the case where learning error is entirely eliminated.
@@
  • James et al. (2013) James, G., D. Witten, T. Hastie, and R. Tibshirani (2013). An Introduction to Statistical Learning Volume 112. Springer.
  • Pedro (2000) Domingos, P. (2000). A Unified Bias-Variance Decomposition and its Applications. In 17th International Conference on Machine Learning, pp.  231–238.